40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 12 11:04:27 2021, Paco, Update, General, PSL shutter was closed this morning 
    Reply  Thu Aug 12 14:59:25 2021, Koji, Update, General, PSL shutter was closed this morning Screen_Shot_2021-08-12_at_14.58.59.png
       Reply  Thu Aug 12 20:52:04 2021, Koji, Update, General, PSL shutter was closed this morning Screen_Shot_2021-08-12_at_20.51.19.pngScreen_Shot_2021-08-12_at_20.51.34.png
Message ID: 16279     Entry time: Thu Aug 12 20:52:04 2021     In reply to: 16278
Author: Koji 
Type: Update 
Category: General 
Subject: PSL shutter was closed this morning 

I did a bit more investigation on this.

- I checked P1~P4, PTP2/3, N2, TP2, TP3. But found only P1a and P2 were affected.

- Looking at the min/mean/max of P1a and P2 (Attachment 1), the signal had a large fluctuation. It is impossible to have P1a from 0.004 to 0 instantaneously.

- Looking at the raw data of P1a and P2 (Attachment 2), the value was not steadily large. Instead it looks like fluctuating noise.

So my conclusion is that because of an unknown reason, an unknown noise coupled only into P1a and P2 and tripped the PSL shutter. I still don't know the status of the mail alert.

Attachment 1: Screen_Shot_2021-08-12_at_20.51.19.png  353 kB  Uploaded Thu Aug 12 21:54:06 2021  | Hide | Hide all
Screen_Shot_2021-08-12_at_20.51.19.png
Attachment 2: Screen_Shot_2021-08-12_at_20.51.34.png  408 kB  Uploaded Thu Aug 12 21:54:06 2021  | Hide | Hide all
Screen_Shot_2021-08-12_at_20.51.34.png
ELOG V3.1.3-